OpenStudy (anonymous):

Can someone explain this to me? Will medal. A system of equations is shown below: 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this.

OpenStudy (anonymous):

@Nnesha @SolomonZelman

OpenStudy (chosenmatt):

still need help? :)

OpenStudy (anonymous):

yes!

OpenStudy (solomonzelman):

I would do elimination.

OpenStudy (anonymous):

hey @bloofoffiction do u know how to use the substitution method?

OpenStudy (anonymous):

I think so

OpenStudy (solomonzelman):

\(\large\color{blue}{8x + 5y = 9 \\ 3x + 2y = 4}\) \(\large\color{blue}{16x + 10y = 18 \\ -15x -10y = -20}\)

OpenStudy (solomonzelman):

I multiplied first equation times 2, and the second times -5. Can you add the equations now?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

so what do you get, after adding?

OpenStudy (anonymous):

u can do that too

OpenStudy (anonymous):

x+x=18 ?

OpenStudy (anonymous):

or 2x=18

OpenStudy (solomonzelman):

well, this is what we are trying to do, \(\large\color{black}{\color{red}{16x} ~~~~~~~~~\color{green}{+ 10y }~~~~=~~~~~~~\color{blue}{ 18} \\\color{red}{-15x}~~~~~\color{green}{- 10y } ~~~~= ~~~~\color{blue}{ -20} }\)

OpenStudy (solomonzelman):

We are adding/subtracting everything separately. 1) the Axes. 2) the whYs. 3) the Constants. (on the right hand side of each equation).

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

so what do you get?

OpenStudy (anonymous):

x=-2 ?

OpenStudy (anonymous):

I'm terrible at algebra

OpenStudy (solomonzelman):

Yes, there you. And the whYs, cancels, as you already figured;)

OpenStudy (solomonzelman):

I meant there you go.

OpenStudy (solomonzelman):

\(\large\color{black}{\color{red}{16x} ~~~~~~~~~\color{green}{+ 10y }~~~~=~~~~~~~\color{blue}{ 18} \\\color{red}{-15x}~~~~~\color{green}{- 10y } ~~~~= ~~~~\color{blue}{ -20} }\) gives as, as you already said, and were correct, \(\large\color{black}{x=-2 }\)

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

So, once you found the x, you can plug it back for to find the y, into any of the original equations. If you plug it into both of the equatins, and find te same result for y, this will then not only solve for x, but also verify that the solution is correct.

OpenStudy (solomonzelman):

reminding, the original equations are: \(\large\color{black}{\color{red}{8x} ~~~~~\color{green}{+ 5y }~~~~=~~~~\color{blue}{ 9} \\\color{red}{3x}~~~~~\color{green}{+2y } ~~~~= ~~~~\color{blue}{ 4} }\)

OpenStudy (anonymous):

\[8(-2)+5y=9\]

OpenStudy (anonymous):

\[3(-2)+2y=4\]

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

what do you get in each equation?

OpenStudy (anonymous):

\[16+5y=9\]

OpenStudy (solomonzelman):

what is that for?

OpenStudy (anonymous):

My self mostly

OpenStudy (anonymous):

I think I've done something wrong because I got y=-1.4

OpenStudy (solomonzelman):

yes, you got something wrong, lol... let s fix that:)

OpenStudy (anonymous):

I would subtract 16 from 9...

OpenStudy (solomonzelman):

\(\large\color{black}{ 8(-2) + 5y = 9 \\ 3(-2) + 2y = 4 }\) \(\large\color{black}{ -16 + 5y = 9 \\ -6 + 2y = 4 }\)

OpenStudy (solomonzelman):

you are adding, because the 16 and the 6 are negative. right?

OpenStudy (anonymous):

oops my bad

OpenStudy (anonymous):

so y = 5 ?

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

and the equivalent systems would be \[16x+10y=18\] and \[-15x-10y=-20\] ?

OpenStudy (solomonzelman):

the first equation si riht, let me check the other one.

OpenStudy (solomonzelman):

yes, and the second one is correct as well.

OpenStudy (anonymous):

ok. Thank you so much! You're the best.

OpenStudy (solomonzelman):

You can use my thing for this. `\(\large\color{black}{ 8x + 5y = 9 \\3x + 2y = 4 }\)` try to copy paste what is in gray, please.

OpenStudy (anonymous):

\(\large\color{black}{ 8x + 5y = 9 \\3x + 2y = 4 }\)

OpenStudy (anonymous):

can I ask a quick question?

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

Why did you multiply the second equation by -5 not positive 5. I need to show the steps and I would like to be able to explain it.

OpenStudy (solomonzelman):

I could have multiplied times 5 and subtracted the equations, but then you confuse is it like \(\large\color{black}{ a-b }\) or \(\large\color{black}{ b-a }\), but in when we are adding, it would matter which way we are going, \(\large\color{black}{ a+(-b) }\) or \(\large\color{black}{ (-b)+a }\)

OpenStudy (solomonzelman):

I mean it wouldn't matter, for the last hrase.

OpenStudy (anonymous):

ok thanks

OpenStudy (solomonzelman):

Anytime;)

OpenStudy (solomonzelman):

Also, if you want to type the \ symbol. 1) Click and hold ALT. 2) Click 2 9 on the number pad on the left corner of the keyboard. 3) Release the ALT key. you get, \ .

OpenStudy (solomonzelman):

So if you are typing in a parenthesis latex, ` \(\large\color{black}{ }\) ` you can do it without copy pasting.

OpenStudy (anonymous):

ok I'll keep that in mind

OpenStudy (solomonzelman):

`\(\large\color{black}{ }\)` this gives you an advantage of typing inside the text, like " \(\large\color{black}{ a+(-b) }\) is the same thing as \(\large\color{black}{ (-b) +a }\)."

OpenStudy (solomonzelman):

yes, don't want to be too annoying... you welcome, once again:)